OpenStudy (anonymous):

#1 What is the vertex of y = |x + 1| ? a. (–1, 0) b. (0, –1) c. (0, 1) d. (1, 0)

OpenStudy (anonymous):

c. (0, 1)

OpenStudy (anonymous):

X + 1 = 0 X = 1 y = |x + 1| y = |1 + 1| y = |-1 + 1| y = 0 is the vertex 1, 0

OpenStudy (anonymous):

@greysica how come its not the other way around?

OpenStudy (anonymous):

if x = 0 then y= |0| +|1|= 1

OpenStudy (anonymous):

oh, okay

OpenStudy (anonymous):

@ganeshie8 is this correct?

OpenStudy (anonymous):

if x = 1 then y= |1|+|1|= 2

OpenStudy (anonymous):

ohhh, i see now.

ganeshie8 (ganeshie8):

to find x-coordinate of vertex, equate the the thing in mod to 0 x+1 = 0 x = -1

ganeshie8 (ganeshie8):

to find y-coordinate of vertex, just plugin x = -1 in the equation y = |x-1| = |1-1| = 0 so the vertex is (-1, 0)

ganeshie8 (ganeshie8):

ur steps in ur first reply are correct. only u made a mistake during arithmetic

OpenStudy (anonymous):

if x= -1 then y=|-1|+|1|= 2

OpenStudy (anonymous):

yeah i see where i went wrong.

ganeshie8 (ganeshie8):

even i made a typo lol : my above reply should read as : to find y-coordinate of vertex, just plugin x = -1 in the equation y = |x+1| = |-1+1| = 0 so the vertex is (-1, 0)

ganeshie8 (ganeshie8):

hope it makes more sense now :P

OpenStudy (anonymous):

okay , I'm getting confuse

OpenStudy (anonymous):

look at what he did. it makes sense.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Hey, thanks a lot @maheshmeghwal9

OpenStudy (maheshmeghwal9):

http://www.wolframalpha.com/input/?i=y%3D |x%2B1|

OpenStudy (maheshmeghwal9):

sorry i m trying but i can't give the graph to y=x+1 Plz u can try actually my net problem:D

OpenStudy (anonymous):

I did, it works!

OpenStudy (maheshmeghwal9):

:)

OpenStudy (anonymous):

(: